OpenStudy (anonymous):

Avery and Collin were trying to challenge each other with equations for sequences. Avery was looking at an explicit equation that Collin wrote. Help (Html5)⇔Help (Java) t(n) = 4.5n – 8 Write the first 4 terms for the sequence.

OpenStudy (anonymous):

ignore the html java thing

OpenStudy (anonymous):

The answer is that they are losers. What people would do math for fun?

OpenStudy (anonymous):

The real answer however, idk

OpenStudy (anonymous):

no, please help me for real

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

who's that?

OpenStudy (anonymous):

He is a person that may be able to help you with your question.

OpenStudy (anonymous):

@ganeshie8 plz help me

ganeshie8 (ganeshie8):

\(t(n) = 4.5n – 8 \) for first term, plugin \(n = 1\)

ganeshie8 (ganeshie8):

first term = \(t(1) = 4.5(1) – 8 = 4.5 - 8 = ? \)

OpenStudy (anonymous):

the 1st term is -3.5 but i cant find the 2nd term

ganeshie8 (ganeshie8):

on the same veins, for second term we plugin \(n=2\)

OpenStudy (anonymous):

is it 1 or -1?

ganeshie8 (ganeshie8):

second term = \(t(2) = 4.5(2) – 8 = 9- 8 = 1 \)

OpenStudy (anonymous):

thanx, my answer key said -1 but i thought it was 1

ganeshie8 (ganeshie8):

np :) you're right !

OpenStudy (anonymous):

I knew he could help :)

OpenStudy (beccaboo333):

Well duh, he's like amazing >v< xD my hunnybear is a math WIZARD!!!!

ganeshie8 (ganeshie8):

ahh so much praise lol xD

OpenStudy (anonymous):

lol hunnybear?

OpenStudy (beccaboo333):

lol casually leaves posts~